# Magneto-optical trap loading with an effusive oven in    a large optical access experiment

### v1

@@ A. Radiation pressure force
${I}^{\pm}_{\mathsf{x}^{\prime};\mathbbm{1}}(\textbf{r})$ $=C\left(y_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),z_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),h_{\mathsf{x};\mathbbm{1}}^{\pm}\right)\times{I}^{\pm}_{\mathsf{x},0;\mathbbm{1}}e^{-2\frac{\left(y_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}+\left(z_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}}{\left(w^{\pm}_{\mathsf{x},0;\mathbbm{1}}\right)^{2}}}$ (A9)
${I}^{\pm}_{\mathsf{y}^{\prime};\mathbbm{1}}(\textbf{r})$ $=C\left(x_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),z_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),h_{\mathsf{y};\mathbbm{1}}^{\pm}\right)\times{I}^{\pm}_{\mathsf{y},0;\mathbbm{1}}e^{-2\frac{\left(x_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}+\left(z_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}}{\left(w^{\pm}_{\mathsf{y},0;\mathbbm{1}}\right)^{2}}}$
${I}^{\pm}_{\mathsf{z}^{\prime};\mathbbm{1}}(\textbf{r})$ $=C\left(x_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),y_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),h_{\mathsf{z};\mathbbm{1}}^{\pm}\right)\times{I}^{\pm}_{\mathsf{z},0;\mathbbm{1}}e^{-2\frac{\left(x_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}+\left(y_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}}{\left(w^{\pm}_{\mathsf{z},0;\mathbbm{1}}\right)^{2}}}$
@@
$y_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})&=\mp\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})x\pm\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})y\pm\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})z\\ z_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})&=\pm\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})x\pm\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})z\\ x_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})&=\pm\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})x\pm\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})y\pm\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})z\\ z_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})&=\mp\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})y\pm\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})z\\ x_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})&=\pm\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})x\pm\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})y\mp\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})z\\ y_{\mathsf{z};\mathbbm{1}^{\prime}}^{\pm}(\textbf{r})&=\pm\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})y\pm\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})z\end{aligned}$ (A10)
@@
are the rotated coordinates, with the offsets supressed for brevity (they can be reintroduced by performing appropriate subtractions to $x,y,z$); and $C(\cdot,:,h_{\alpha;\mathbbm{1}}^{\pm})$ is the cylinder function with $h_{\alpha;\mathbbm{1}}^{\pm}$ being its cut-off ($C=1$ for $\sqrt{(\cdot)^{2}+(:)^{2}}\leq h_{\alpha;\mathbbm{1}}^{\pm}$, and 00 otherwise). The truncation is naturally considered, as the beams pass through the optics apertures limiting their size.
@@
The final quantity in Eq. A3, $\sigma_{\alpha^{\prime},q;\mathbbm{1}}^{\pm}$, is the corresponding scattering cross-section for a single 2-level atomic transition:
@@
$\hskip 22.0pt\sigma^{\pm}_{\alpha^{\prime},q;\mathbbm{1}}(\textbf{r},\textbf{v})=\frac{\sigma_{0}}{1+\frac{I_{tot,q}(\textbf{r})}{I_{sat}}+4\frac{(\Delta_{\mathbbm{1}}-\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{v}-\mu_{q}(\textbf{r}))^{2}}{\Gamma^{2}}}$ (A11)
@@
where $\sigma_{0}=6\pi/k_{L}^{2}$ is the resonant scattering cross section; the detuning $\Delta_{\mathbbm{1}}=\Delta_{\textsf{3M}}$, $\Delta_{\textsf{2M}}$, $\Delta_{\textsf{2O}}$, or $\Delta_{\textsf{ZS}}$ is that of the 3D MOT, 2D MOT, 2D OM, or ZS beams, respectively; $\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{v}$ is the Doppler shift for a positive or negative ($\pm$) beam; and $\mu_{q}(\textbf{r})=q\mu B(\textbf{r})$ is the Zeeman shift for the $m_{1}=-1,0,+1$ level (where, respectively, $q=-,0,+$), with $\mu$ being the gyromagnetic ratio (the particular Zeeman shift is due to the quantization axis being chosen to be along the direction of B);
[…]
@@
is the total beam intensity that a single 2-level transition receives, which is seen to be a sum of the corresponding total beam intensities of the 3D MOT, 2D MOT, 2D OM, and ZS, respectively. For the 3D MOT, for instance, one has $I_{tot,q;\textsf{3M}}=\sum_{\alpha^{\prime}}p_{\alpha^{\prime},q;\textsf{3M}}^{+,Q}(\textbf{r}){I}^{+}_{\alpha^{\prime};\textsf{3M}}(\textbf{r})+p_{\alpha^{\prime},q;\textsf{3M}}^{-,Q}(\textbf{r}){I}^{-}_{\alpha^{\prime};\textsf{3M}}(\textbf{r})$, where $Q=\sigma^{-}$ for $\alpha^{\prime}=\mathsf{x}^{\prime},\mathsf{y}^{\prime}$ and $Q=\sigma^{+}$ for $\alpha^{\prime}=\mathsf{z}^{\prime}$.
@@ B. Diffusion
The fluctuating part of the trapping force is here introduced via a momentum diffusion coefficient that we write as follows:
$D(\textbf{r},\textbf{v})=\hbar^{2}k^{2}_{L}\frac{\Gamma}{4}\frac{s_{tot}(\textbf{r},\textbf{v})}{1+s_{tot}(\textbf{r},\textbf{v})}$ (A13)
@@
$s_{tot}(\textbf{r},\textbf{v})=\sum_{\mathbbm{1}}s_{\mathbbm{1}}(\textbf{r},\textbf{v})$ (A14)
[…]
$s_{\mathbbm{1}}(\textbf{r},\textbf{v})=\sum_{q=\sigma^{+},\sigma^{-},\pi}s_{q;\mathbbm{1}}(\textbf{r},\textbf{v})$ (A15)
@@
$s_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,Q}(\textbf{r},\textbf{v})=\frac{p_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,Q}(\textbf{r}){I}^{\pm}_{\alpha^{\prime};\mathbbm{1}}(\textbf{r})/I_{sat}}{1+4\frac{(\Delta_{\mathbbm{1}}-\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{v}-\mu_{q}(\textbf{r}))^{2}}{\Gamma^{2}}}$ (A16)
@@
From Eq. (A2) in Ref. [48] for the steady-state atom number in the excited blue cooling level $5s5p$ ${}^{1}P_{1}$, we can
Figure B1: Color online. Diagrams showing the normalized 3D MOT atom numbers obtained using our analytical model describing the 3D MOT loading using the moving molasses technique discussed in the main text (Sec. III). The left diagram uses the full Eq. 1 of our model, while the right one excludes the loss to the dark state $5s5p\,^{3}\!P_{0}$ ($\mathcal{F}_{0}\tau_{\ominus}=1$ is assumed in Eq. 1). The left dashed line and the cross are as in Fig. 2, while the right dashed line indicates the boundary where the sign of the confinement is flipped. Note that the diagrams span over a larger parameter space than in Fig. 2.
obtain the following expression for the mean atom-lifetime:
$\tau=\frac{1}{\Gamma_{23}^{\prime}}\frac{\Gamma_{34}+\Gamma_{36}}{\Gamma_{34}}\mathcal{E}\quad\quad\quad\quad\quad\;\;\,\text{(A17)}\end{aligned}$
@@
$\Gamma_{23}^{\prime}=\frac{\Gamma_{23}}{2}\frac{s_{tot}(\textbf{r},\textbf{v})}{1+s_{tot}(\textbf{r},\textbf{v})}\quad\quad\quad\quad\quad\text{(A18)}\end{aligned}$
is the photon scattering rate for the transition $5s5p\,^{1}\!P_{1}\rightarrow 5s4d\,^{1}\!D_{2}$ due to the blue cycling with the saturation $s_{tot}$ (Eq. A14), which we note introduces  dependency  on the local intensity, magnetic field, and velocity. This substitution is motivated by Eq. (1) in Ref. [51], where one can identify $A_{{}^{1}\!P_{1}\rightarrow^{1}\!D_{2}}=\Gamma_{23}=3.9\times 10^{3}$ Hz as the corresponding transition linewidth, and $B_{{}^{1}\!D_{2}\rightarrow^{3}\!P_{2}}=\left(\frac{\Gamma_{34}+\Gamma_{36}}{\Gamma_{34}}\right)^{-1}=0.33$ as the branching ratio for the shelving transition $5s4d\,^{1}\!D_{2}\rightarrow 5s5p\,^{3}\!P_{2}$ (with the corresponding linewidth $\Gamma_{34}=6.6\times 10^{2}$ Hz, and $\Gamma_{36}=1.34\times 10^{3}$ Hz being the linewidth for $5s4d\,^{1}\!D_{2}\rightarrow 5s5p\,^{3}\!P_{1}$).
@@
We note that, as mentioned in the main text, Sec. II, the lifetime of a superparticle in the simulations is determined by multiplying $\tau$ by a number predrawn from a unit exponential distribution.
Appendix B
@@
Attached to this appendix is Fig. B1 that shows two diagrams calculated using our analytical model (Sec. III.A), in order to support our main observations (Sec. III.B).

### v2

@@ A. Radiation pressure force
$\hskip 115.0pt\begin{aligned} \hskip-110.0pt{I}^{\pm}_{\mathsf{x}^{\prime};\mathbbm{1}}(\textbf{r})&=C\left[y_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),z_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),h_{\mathsf{x};\mathbbm{1}}^{\pm}\right]\times{I}^{\pm}_{\mathsf{x},0;\mathbbm{1}}e^{-2\frac{\left(y_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}+\left(z_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}}{\left(w^{\pm}_{\mathsf{x},0;\mathbbm{1}}\right)^{2}}}\\ \hskip-110.0pt{I}^{\pm}_{\mathsf{y}^{\prime};\mathbbm{1}}(\textbf{r})&=C\left[x_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),z_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),h_{\mathsf{y};\mathbbm{1}}^{\pm}\right]\times{I}^{\pm}_{\mathsf{y},0;\mathbbm{1}}e^{-2\frac{\left(x_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}+\left(z_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}}{\left(w^{\pm}_{\mathsf{y},0;\mathbbm{1}}\right)^{2}}}\\ \hskip-110.0pt{I}^{\pm}_{\mathsf{z}^{\prime};\mathbbm{1}}(\textbf{r})&=C\left[x_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),y_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r}),h_{\mathsf{z};\mathbbm{1}}^{\pm}\right]\times{I}^{\pm}_{\mathsf{z},0;\mathbbm{1}}e^{-2\frac{\left(x_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}+\left(y_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})\right)^{2}}{\left(w^{\pm}_{\mathsf{z},0;\mathbbm{1}}\right)^{2}}}\end{aligned}$ (A9)
@@
$\hskip 20.0pt\begin{aligned} \hskip-113.0pty_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})&=\mp\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})x\pm\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})y\pm\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})z\\ \hskip-113.0ptz_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})&=\pm\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})x\pm\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})z\\ \hskip-113.0ptx_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})&=\pm\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})x\pm\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})y\pm\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}})z\\ \hskip-113.0ptz_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})&=\mp\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})y\pm\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})z\\ \hskip-113.0ptx_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm}(\textbf{r})&=\pm\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})x\pm\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})y\mp\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})z\\ \hskip-113.0pty_{\mathsf{z};\mathbbm{1}^{\prime}}^{\pm}(\textbf{r})&=\pm\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})y\pm\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})z\end{aligned}$ (A10)
@@
are the rotated coordinates, with the offsets supressed for brevity (they can be reintroduced by performing appropriate subtractions to $x,y,z$); and $C[\cdot,:,h_{\alpha;\mathbbm{1}}^{\pm}]$ is the cylinder function with $h_{\alpha;\mathbbm{1}}^{\pm}$ being its cut off [$C=1$ for $\sqrt{(\cdot)^{2}+(:)^{2}}\leq h_{\alpha;\mathbbm{1}}^{\pm}$, and 0 otherwise]. The truncation is naturally considered, as the beams pass through the optics apertures limiting their size.
@@
The final quantity in Eq. A3, $\sigma_{\alpha^{\prime},q;\mathbbm{1}}^{\pm}$, is the corresponding scattering cross section for a single two-level atomic transition:
@@
$\hskip-90.0pt\sigma^{\pm}_{\alpha^{\prime},q;\mathbbm{1}}(\textbf{r},\textbf{v})=\frac{\sigma_{0}}{1+\frac{I_{tot,q}(\textbf{r})}{I_{sat}}+4\frac{(\Delta_{\mathbbm{1}}-\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{v}-\mu_{q}(\textbf{r}))^{2}}{\Gamma^{2}}}$ (A11)
@@
where $\sigma_{0}=6\pi/k_{L}^{2}$ is the resonant scattering cross section; the detuning $\Delta_{\mathbbm{1}}=\Delta_{\textsf{3M}}$, $\Delta_{\textsf{2M}}$, $\Delta_{\textsf{2O}}$, or $\Delta_{\textsf{ZS}}$ is that of the 3D MOT, 2D MOT, 2D OM, or ZS beams, respectively; $\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{v}$ is the Doppler shift for a positive or negative ($\pm$) beam; and $\mu_{q}(\textbf{r})=q\mu B(\textbf{r})$ is the Zeeman shift for the $m_{1}=-1,0,+1$ level (where, respectively, q=,0,+q=-,0,+), with $\mu$ being the gyromagnetic ratio (the particular Zeeman shift is due to the quantization axis being chosen to be along the direction of B);
$\hskip-92.0ptI_{tot,q}(\textbf{r})=\sum_{\mathbbm{1}}I_{tot,q;\mathbbm{1}}(\textbf{r},\textbf{v})$ (A12)
@@
is the total beam intensity that a single two-level transition receives, which is seen to be a sum of the corresponding total beam intensities of the 3D MOT, 2D MOT, 2D OM, and ZS, respectively. For the 3D MOT, for instance, one has $I_{tot,q;\textsf{3M}}=\sum_{\alpha^{\prime}}p_{\alpha^{\prime},q;\textsf{3M}}^{+,Q}(\textbf{r}){I}^{+}_{\alpha^{\prime};\textsf{3M}}(\textbf{r})+p_{\alpha^{\prime},q;\textsf{3M}}^{-,Q}(\textbf{r}){I}^{-}_{\alpha^{\prime};\textsf{3M}}(\textbf{r})$, where $Q=\sigma^{-}$ for $\alpha^{\prime}=\mathsf{x}^{\prime},\mathsf{y}^{\prime}$ and $Q=\sigma^{+}$ for $\alpha^{\prime}=\mathsf{z}^{\prime}$.
@@ B. Diffusion
The fluctuating part of the trapping force is introduced here via a momentum diffusion coefficient that we write as follows:
$\hskip 57.0pt\begin{aligned} D(\textbf{r},\textbf{v})=\hbar^{2}k^{2}_{L}\frac{\Gamma}{4}\frac{s_{tot}(\textbf{r},\textbf{v})}{1+s_{tot}(\textbf{r},\textbf{v})}\quad\quad\quad\quad\text{(A13)}\end{aligned}$
@@
$\hskip 75.0pt\begin{aligned} s_{tot}(\textbf{r},\textbf{v})=\sum_{\mathbbm{1}}s_{\mathbbm{1}}(\textbf{r},\textbf{v})\quad\quad\quad\quad\quad\,\,\text{(A14)}\end{aligned}$
is the total saturation parameter that is a sum of the individual total saturation parameters $s_{\mathbbm{1}}$, where $\mathbbm{1}=\textsf{3M},\textsf{2M},\textsf{2O}$, or ZS (respectively referring to the 3D MOT, 2D MOT, 2D OM, or ZS). We write $s_{\mathbbm{1}}$ as the sum of the saturation parameters $s_{q;\mathbbm{1}}$ for the atom’s two-level transitions that are driven by, respectively, $\sigma^{-}$, $\pi$, $\sigma^{+}$ polarized light:
$\hskip 65.0pt\begin{aligned} s_{\mathbbm{1}}(\textbf{r},\textbf{v})=\sum_{q=\sigma^{+},\sigma^{-},\pi}s_{q;\mathbbm{1}}(\textbf{r},\textbf{v})\quad\quad\quad\;\;\text{(A15)}\end{aligned}$
@@
$\hskip 45.0pt\begin{aligned} s_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,Q}(\textbf{r},\textbf{v})=\frac{p_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,Q}(\textbf{r}){I}^{\pm}_{\alpha^{\prime};\mathbbm{1}}(\textbf{r})/I_{sat}}{1+4\frac{(\Delta_{\mathbbm{1}}-\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{v}-\mu_{q}(\textbf{r}))^{2}}{\Gamma^{2}}}\quad\;\;\text{(A16)}\end{aligned}$
@@
From Eq. (A2) in Ref. [50] for the steady-state atom number in the excited blue cooling level $5s5p$ $^{1}P_{1}$, we can obtain the following expression for the mean atom-lifetime:
$\hskip 90.0pt\begin{aligned} \tilde{\tau}=\frac{1}{\Gamma_{23}^{\prime}}\frac{\Gamma_{34}+\Gamma_{36}}{\Gamma_{34}}\mathcal{E}\quad\quad\quad\quad\quad\,\,\,\,\text{(B1)}\end{aligned}$
@@
$\hskip 80.0pt\begin{aligned} \Gamma_{23}^{\prime}=\frac{\Gamma_{23}}{2}\frac{s_{tot}(\textbf{r},\textbf{v})}{1+s_{tot}(\textbf{r},\textbf{v})}\quad\quad\quad\quad\quad\text{(B2)}\end{aligned}$
is the photon scattering rate for the transition $5s5p\,^{1}\!P_{1}\rightarrow 5s4d\,^{1}\!D_{2}$ due to the blue cycling with the saturation $s_{tot}$ (Eq. A14), which we note introduces  dependency  on the
Figure 7: Diagrams showing the normalized 3D MOT atom numbers obtained using our analytical model describing the 3D MOT loading using the moving molasses technique discussed in the main text (Sec. III). The left diagram uses the full Eq. 3 of our model, while the right one excludes the loss to the dark state $5s5p\,^{3}\!P_{0}$ ($\Gamma_{\text{loss}}$ is held at a fixed positive value in Eq. 3). The left dashed line and the cross are as in Fig. 4, while the right dashed line indicates the boundary where the sign of the confinement is flipped. Note that the diagrams span over a larger parameter space than in Fig. 4.
local intensity, magnetic field, and velocity. This substitution is motivated by Eq. (1) in Ref. [53], where one can identify $A_{{}^{1}\!P_{1}\rightarrow^{1}\!D_{2}}=\Gamma_{23}=3.9\times 10^{3}$ Hz as the corresponding transition linewidth, and $B_{{}^{1}\!D_{2}\rightarrow^{3}\!P_{2}}=\left(\frac{\Gamma_{34}+\Gamma_{36}}{\Gamma_{34}}\right)^{-1}=0.33$ as the branching ratio for the shelving transition $5s4d\,^{1}\!D_{2}\rightarrow 5s5p\,^{3}\!P_{2}$ (with the corresponding linewidth $\Gamma_{34}=6.6\times 10^{2}$ Hz, and $\Gamma_{36}=1.34\times 10^{3}$ Hz being the linewidth for $5s4d\,^{1}\!D_{2}\rightarrow 5s5p\,^{3}\!P_{1}$).
@@
We note that as mentioned in Sec. II, the lifetime of a superparticle in the simulations is determined by multiplying $\tilde{\tau}$ by a number that is predrawn from a unit exponential distribution.
Appendix C
@@
Figure 7 shows two diagrams calculated using our analytical model (Sec. III.A), in order to support our main observations (Sec. III.B). See the figure text for details.
Magneto-optical-trap loading in a large optical-access experiment
               (Supplemental Material)
We here demonstrate our simulation tool’s ability to improve the performance of an effusive oven experiment. We discuss how the 2D MOT loading can be enhanced by introducing changes to the 2D OM and the ZS. Also, we numerically verify the optimal angles for the radial beams in the 3D MOT.
Two important ways to manipulate the 2D OM performance include optimizing the beam power and the beam size (aperture and waist) [37]. The former is, however, resource intensive, considering the gains diminish rapidly due to the atom saturation while other parts of the setup (e.g., the 2D and 3D MOTs) may benefit more from the available laser power. In Fig. S.1, we study the effects of increased beam power for different waist sizes at fixed aperture sizes. The data points connected with a straight line are obtained for the original aperture and beam waist (Tab. 1 case 4), while the data points connected with other kinds of lines are for a $4$ times larger aperture (requiring two-inch optics) and different beam waists (dashed for $1$ times the original, dash-dotted for $2$, and dotted for $3$). We observe first that enlarging the aperture can result in an increased 2D MOT loading rate, depending on both the beam waist and the 2D OM power. An increase in this rate is a consequence of the atoms being exposed longer to the transverse cooling, provided the intensity is sizeable throughout the cooling region with the chosen beam waist. Indeed, as observed for lower intensities (below 30 mW), extending the waist can either be advantageous (dashed and dash-dotted lines) due to effectively improved cooling, or detrimental (dotted line) due to effectively diminished cooling. For larger intensities, on the other hand, the saturation effect is present, such that wider waists result in the extension of a region of efficient cooling. In this case, it is particularly important to choose a waist that is not too narrow, so this effect can be taken full  advantage  of  (observe the  dotted  line
crossing the dashed one above 30 mW). Overall, the observations here agree with intuitive expectations, further validating the numerical model.
Figure S.1: Numerical data showing the 2D MOT loading rate versus the 2D OM beam power (individual). Different aperture radii $a$ and beam waist radii $w$ of the 2D OM are considered. The aperature size $a\times 1$” and the waist size $w\times 1$” are from the original simulations, where $6$ mW of 2D OM beam power is also used (Tab. 1 case 4). The loading normalization is done with respect to this case.
Our ZS modeling (Eq. A8) has been derived with a purpose of tailoring a magnetic field to a given experimental environment, and we outline here how this can be achieved. With the initial oven conditions, ZS beam and 2D MOT parameters specified, one employs an algorithm that iteratively adjusts the slopes of the linear segments composing the field profile, with the aim of increasing the probability density of the atom speeds within the capture-speed range of the 2D MOT. To remove spurious fluctuations, each iteration step uses the same initial
atom position and speed distributions with the diffusion (stemming from the radiation pressure; App. A1.B) turned off. In Fig. S.2, we display the initial speed profile (light gray), and the final speed profile at the 2D MOT location (dark gray), which can be achieved in the iteration process. A concentrated probability density is seen to occur close to 0 m/s for the final profile (see the inset), corresponding to the atoms that have been captured by the 2D MOT. The employed field profile stems from our experiment where it more than quadruples the 2D MOT loading rate (see Tab. I), but it has not been optimized using the numerical approach outlined above. We expect it to yield accurate optimization results, given the closely matching numerical and experimental 2D MOT loading rates obtained with help of the segmentation, and possibly achieve a 20-fold enhancement with appropriate ZS stage modifications [37, 40]. We note that our numerical approach is similar to the one extensively tested in Ref. [40], with the main difference there being that the motion is in 1D and the radial expansion of the oven flux is approximated to yield an effective 3D simulation. We note that although the lower (actual) dimensionality will result in greater computational efficiency, this difference may lead to decreased accuracy compared to our approach.
Finally, we verify for the 3D MOT side beams their optimal angles with respect to the radial axis perpendicular to the objectives axis. In Fig. S.3, we display the numerical results for the 3D MOT atom numbers versus the beams angle (mirrored). As can be seen, the $25\,^{\circ}$ angle yields the highest number given our experimental constrains (dashed line limit). While larger angles can result in even higher numbers (see right of the dashed line), the beams may be clipped by our objectives [refer to the inset in Fig. 2(a)]. The optimum angle is found to be greater, at $30\,^{\circ}$, which can be explained to be a combination of a large repumping light volume in the science region (thus bringing many shelved atoms back into the cooling cycle) and the radial trapping forces being closer to perpendicular (thus resulting in more balanced forces). The resulting increase in the atom number is, however, marginal (by ${\sim}25\,\%$), and it is relatively robust against changes within a wide range of angles (from $20\,^{\circ}$ to $45\,^{\circ}$). Note that the findings here are not universal due to their dependence on the cell geometry. Implicitly, however, we are made aware that the optimum loading is in general not  achieved  at  the  standard $90\,^{\circ}$ angle  between  beam
axes (corresponding to $45\,^{\circ}$ in Fig. S.3).
Figure S.2: Numerical data showing atom speed histograms before the Zeeman slowing (light gray) and at the 2D MOT location after the slowing (dark gray). The solid line displays the Maxwell-Boltzmann distribution for $440\,^{\circ}$C (as used in the original simulations; see Sec. II.A). The inset zooms in on the capturable speeds of the 2D MOT, showing a concentrated probability density near 0 m/s due to the trapping.
Figure S.3: Numerical data showing the 3D MOT atom number versus the angle of the side beams mirrored with respect to the radial axis perpendicular to the objectives axis [refer to the inset in Fig. 2(a)]. The vertical dashed line indicates the limit beyond which our objectives may clip the beams. The normalization is done with respect to the point at the $25\,^{\circ}$ angle, as used in the experiment (see Sec. II.D). The remaining simulation parameters are the same as in Fig. 4, where the greatest atom number is obtained.